OpenStudy (anonymous):

write the equation of the line passing through the points (-3,2) (1,-2)....please explain the steps =]

OpenStudy (saifoo.khan):

mimi wannna try?

OpenStudy (saifoo.khan):

Hello Meeeeera.

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

can you please help me ...? =]

OpenStudy (saifoo.khan):

Sure, mini is sleepping.

OpenStudy (mimi_x3):

y=-x - 2

OpenStudy (saifoo.khan):

sleeping*

OpenStudy (mimi_x3):

Pls correct the other typo as well (:

OpenStudy (saifoo.khan):

BLIND she wrote WITH STEPS.

OpenStudy (mimi_x3):

Woops, I woke up i didnt read it properly

OpenStudy (saifoo.khan):

OPEN UP YOUR EYES! O_O

OpenStudy (anonymous):

wait so how do i solve this?

OpenStudy (mimi_x3):

Gradient formula y^2-y^1/x^2-x^1 = -1 y=mx +b 2 = -1(-3)+b 2=3+ b b =-1

OpenStudy (saifoo.khan):

RIght.

OpenStudy (mimi_x3):

Obviously its right :P

OpenStudy (saifoo.khan):

meera are u getting it?

OpenStudy (mimi_x3):

It's common sese she should get it :P

OpenStudy (saifoo.khan):

sense*

OpenStudy (mimi_x3):

Only I have the honour to correct my typo :P

OpenStudy (saifoo.khan):

Where is that rule written?

OpenStudy (mimi_x3):

In my vocab :P

OpenStudy (saifoo.khan):

i nvr read it. Show me. :p

OpenStudy (anonymous):

thnx saifoo... i get it ♥ =]

OpenStudy (saifoo.khan):

No Problem! <3

OpenStudy (mimi_x3):

Um, btw i was the one who showed the working out NOT Saifoo :P

OpenStudy (anonymous):

lol thank you =]

OpenStudy (mimi_x3):

lols

OpenStudy (saifoo.khan):

HAHA @ mimi.